OpenStudy (anonymous):

How do I figure out: 35% of what number is 17.5?

OpenStudy (anonymous):

can anyone help me with proportions? please?

OpenStudy (anonymous):

35/100 = 17.5/x

OpenStudy (anonymous):

take your 17.5 and multiply it by .35

OpenStudy (darthsid):

Here's how to approach these types of problems: Let the number in question be x. Now, you are given that 35% of x is 17.5 This can be written as :

OpenStudy (darthsid):

Sorry, accidentally posted above. 35/100 x = 17.5 Multiply both sides by hundred, 35x = 1750 Now, you can divide both sides by 35 to get the answer.

OpenStudy (anonymous):

Thank you! What about: What number is 23% of 16? How?

OpenStudy (anonymous):

23 percent is 23 of 100. Your proportion should look like: 23 16 --- = --- 100 x Solve by cross multiplying: 23x=16*100

OpenStudy (anonymous):

But then what? is the answer 23x =1600? Don't i need to get x by itself? To solve?